OpenStudy (anonymous):

When 2 less than a number n is increased by 70%, the result is 8.5. Write an equation to represent this statement and find the value of n.

OpenStudy (anonymous):

Let n = the number. Then 2 less then this number would be: n-2 Next we are told that (n-2) is increased by 70%. So we have the original value plus we have to add 70% more to it. So we now have (n-2) + .70(n-2) Next we are told that original value plus the increase of 70% equals 8.5 So we now have this (n-2) + .70(n-2) = 8.5 Now we just have to do a little housekeeping. So we have this: (n-2) ( 1+.70) = 8.5 distributive law (n-2)(1.7) = 8.5 combine like terms I probably wouldn't reduce the equation any more, since you asked to represent the statement. To find n, you will have to keep going. n-2 = 8.5/1.7 n-2=5 n=7